OpenStudy (anonymous):

two numbers total -2, and their difference is 4. Find the two numbers

OpenStudy (anonymous):

x + y = -2 x - y = 4

OpenStudy (anonymous):

Solve one equation for a variable and substitute into the other equation, then find the second variable :) and those are your two numbers

OpenStudy (campbell_st):

let your numbers be a and b then a + b = -2 The sum of the numbers a - b = 4 the difference add the 2 equations to eliminate b 2a = 2 solve for a. When you have a substitute into either equation to find b

OpenStudy (anonymous):